# Multiple SLEs and Dyson Brownian motion: transition density and Green’s function

### v2

@@ 1 Introduction
When $N=1$, we denote by $X_{\mathrm{simple}}(\Omega;x_{1},x_{2})$ the set of continuous simple unparameterized curves in $\Omega$ connecting $x_{1}$ and $x_{2}$ such that they only touch the boundary $\partial\Omega$ in $\{x_{1},x_{2}\}$. We denote by $X(\Omega;x_{1},x_{2})$ the closure of $X_{\mathrm{simple}}(\Omega;x_{1},x_{2})$ in the metric topology of the set $X$ of all planar oriented curves (i.e., continuous mappings from $[0,1]$ to $\mathbb{C}$ modulo reparameterization) equipped with metric
$\mathrm{dist}(\eta,\tilde{\eta}):=\inf_{\varphi,\tilde{\varphi}}\sup_{t\in[0,1]}|\eta(\varphi(t))-\tilde{\eta}(\tilde{\varphi}(t))|,\quad\eta,\tilde{\eta}\in X,$
@@
where the infimum is taken over all increasing homeomorphisms $\varphi,\tilde{\varphi}:[0,1]\mapsto[0,1]$. Chordal SLE in $(\Omega;x_{1},x_{2})$ is a continuous random curve in $\Omega$ connecting $x_{1}$ and $x_{2}$. Note that $\mathrm{SLE}_{\kappa}$ in $(\Omega;x_{1},x_{2})$ is a random curve in $X_{\mathrm{simple}}(\Omega;x_{1},x_{2})$ when $\kappa\leq 4$ and that $\mathrm{SLE}_{\kappa}$ in $(\Omega;x_{1},x_{2})$ is a random curve in $X(\Omega;x_{1},x_{2})$ when $\kappa\in(4,8)$. In general, for $N\geq 2$ and $\alpha=\{\{a_{1},b_{1}\},\ldots,\{a_{N},b_{N}\}\}\in\mathrm{LP}_{N}$, we denote by $X_{\alpha}(\Omega;x_{1},\ldots,x_{2N})$ the set of families $(\gamma_{1},\ldots,\gamma_{N})$ of curves such that, for each $s\in\{1,\ldots,N\}$, we have $\gamma_{s}\in X(\Omega;x_{a_{s}},x_{b_{s}})$ and $\gamma_{s}$ does not disconnect any two points $x_{a},x_{b}$ with $\{a,b\}\in\alpha$ from each other.
@@
Fix $\kappa\in(0,8)$. For $N\geq 1$ and $\alpha\in\mathrm{LP}_{N}$, there exists a unique probability measure on $(\gamma_{1},\ldots,\gamma_{N})\in X_{\alpha}(\Omega;x_{1},\ldots,x_{2N})$ such that, for each $j\in\{1,\ldots,N\}$, the conditional law of $\gamma_{j}$ given $\{\gamma_{1},\ldots,\gamma_{N}\}\setminus\{\gamma_{j}\}$ is the chordal $\mathrm{SLE}_{\kappa}$ connecting $x_{a_{j}}$ and $x_{b_{j}}$ in the connected component of the domain $\Omega\setminus\bigcup_{k\neq j}\gamma_{k}$ having $x_{a_{j}}$ and $x_{b_{j}}$ on its boundary. For the existence and uniqueness of such measures, see [KL07], [PW19, Theorem 1.3] and [BPW21, Theorem 1.2] for $\kappa\in(0,4]$, and [AHSY23, Theorem 1.2], [Zha23, Theorem 4.2] and [FLPW24, Theorem 1.21] for $\kappa\in(4,8)$. We call this measure chordal $N$-$\mathrm{SLE}_{\kappa}$ in polygon $(\Omega;x_{1},\ldots,x_{2N})$ associated to link pattern $\alpha$.
@@
In the companion paper [FWY24], we derived the connection between chordal $N$-$\mathrm{SLE}_{\kappa}$ in the unit disc and Dyson circular ensemble for $\kappa\in(0,4]$. In this article, we will prove corresponding conclusions for chordal $N$-$\mathrm{SLE}_{\kappa}$ and Dyson Brownian motion, and extend the conclusions from $\kappa\in(0,4]$ to $\kappa\in(0,8)$. The main conclusions are summarized as follows. We will parameterize multiple chordal SLEs by a “common time” and derive the transition density of the driving function in Theorem 1.1: it can be given by the transition density of Dyson Brownian motion and Green’s function (Definition 1.2). Using such connection and the asymptotic of the transition density of Dyson Brownian motion, we derive the asymptotic of the probability of a rare event and show that, conditional on this rare event, the driving function of multiple chordal SLEs converges to Dyson Brownian motion (Proposition 1.3). In particular, we take $\kappa=4$ as an example and connect our result to Gaussian free field level lines (Proposition 4.1). In this case, the driving function has an explicit SDE (4.2), Green’s function has an explicit formula (4.4), and the Dyson Brownian motion is of parameter $\beta=2$. Finally, we give one application of our main result: we derive the asymptotic of the probability for chordal $N$-SLE to hit a small neighborhood of a boundary point (Theorem 1.4).
@@
Let us elaborate on how we extend our results from $\kappa\in(0,4]$ to $\kappa\in(0,8)$. To prove the first main conclusion Theorem 1.1, we first relate global multiple SLEs to multiple SLE partition functions “pure partition functions” (see Section 2.1); second, we use these partition functions to construct $2N$-time local martingales (see Section 3.2); and finally, we take the common time in these local martingales and derive the desired transition density. This proof works for all $\kappa\in(0,8)$. To prove the second main conclusion Proposition 1.3 and Theorem 1.4, we need further analysis on pure partition functions in Section 2.2. Such analysis relies on the relation between Coulomb gas integrals and pure partition functions which is only available due to recent development in [FPW24, FLPW24]. Using these tools, we are also able to extend previous result [FWY24, Theorem 1.2] from $\kappa\in(0,4]$ to $\kappa\in(0,8)$, see Remark 1.5.
@@ 1.1 Multiple SLEs and Dyson Brownian motion
Fix $\kappa\in(0,8)$ and $\alpha\in\mathrm{LP}_{N}$.
@@
Fix $\kappa\in(0,8)$ and $\alpha\in\mathrm{LP}_{N}$. For $(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$, we define Green’s function in the boundary for chordal $N$-$\mathrm{SLE}_{\kappa}$ as
@@ 1.2 Estimates on multiple SLEs
Fix $\kappa\in(0,8)$, $\alpha\in\mathrm{LP}_{N}$ and $\boldsymbol{x}\in\mathfrak{X}_{2N}$. We have
@@
Fix $\kappa\in(0,8)$, $\boldsymbol{x}=(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$, $\alpha\in\mathrm{LP}_{N}$ and suppose $(\gamma_{1},\ldots,\gamma_{N})\sim\mathbb{P}_{\alpha}^{(\boldsymbol{x})}$. Then we have [^2] For two functions $F_{1}(\boldsymbol{x};R)$ and $F_{2}(\boldsymbol{x};R)$, we write $F_{1}(\boldsymbol{x};R)\asymp F_{2}(\boldsymbol{x};R)$ as $R/|\boldsymbol{x}|\to\infty$ if there exist constants $C\in(1,\infty)$ and $R_{0}$ such that $C^{-1}\leq\frac{F_{1}(\boldsymbol{x};R)}{F_{2}(\boldsymbol{x};R)}\leq C$ for all $R\geq R_{0}|\boldsymbol{x}|$.
@@
The rest of this article is organized as follows. We give preliminaries on multiple SLEs and Dyson Brownian motion in Section 2. We prove Theorem 1.1 in Section 3. The main ingredients for the proof are $2N$-time local martingales constructed in Section 3.2. Using Theorem 1.1 and the asymptotic of the transition density of Dyson Brownian motion, we complete the proof of Proposition 1.3 in Section 3.3. We prove Proposition 4.1 in Section 4. Finally, we prove Proposition 1.4 in Section 5 using Proposition 1.3. The conclusion in Proposition 1.4 is proved for $\kappa\in(0,8),N=1$ in [Law15] and is proved for $\kappa\in(0,8),N=2$ in [Zha21]. We will prove (1.8) in Section 5 for all $\kappa\in(0,8),N\geq 3,\alpha\in\mathrm{LP}_{N}$.
Remark 1.5.
In the companion paper [FWY24], we derived the asymptotic of the probability for chordal $N$-$\mathrm{SLE}_{\kappa}$ to hit a small neighborhood of an interior point for $\kappa\in(0,4]$. Using recent development in [FPW24, FLPW24], we are able to extend it from $\kappa\in(0,4]$ to $\kappa\in(0,8)$. More precisely, fix $\kappa\in(0,8)$, consider the polygon $(\mathbb{U};x_{1},\ldots,x_{2N})$ where $\mathbb{U}$ is the unit disc. Let $(\gamma_{1},\ldots,\gamma_{N})$ be the chordal $N$-$\mathrm{SLE}_{\kappa}$ in $(\mathbb{U};x_{1},\ldots,x_{2N})$ associated to link pattern $\alpha\in\mathrm{LP}_{N}$ and denote its law by $\mathbb{P}_{\alpha}^{(\mathbb{U};\boldsymbol{x})}$. Then [^3] For two functions $F_{1}(\boldsymbol{x};\epsilon)$ and $F_{2}(\boldsymbol{x};\epsilon)$, we write $F_{1}(\boldsymbol{x};\epsilon)\sim F_{2}(\boldsymbol{x};\epsilon)$ as $\epsilon\to 0$ if the limit $\lim_{\epsilon\to 0}\frac{F_{1}(\boldsymbol{x};\epsilon)}{F_{2}(\boldsymbol{x};\epsilon)}$ exists and takes value in $(0,\infty)$.
$\mathbb{P}_{\alpha}^{(\mathbb{U};\boldsymbol{x})}\left[\gamma_{j}\cap B(0,\epsilon)\neq\emptyset,1\leq j\leq N\right]\sim\tilde{G}_{\alpha}(\boldsymbol{x})\epsilon^{A_{2N}},\quad\text{as }\epsilon\to 0,$ (1.9)
where $A_{2N}$ is whole-plane arm exponent
$A_{2N}=\frac{16N^{2}-(\kappa-4)^{2}}{8\kappa};$
and $\tilde{G}_{\alpha}$ is Green’s function defined in [FWY24, Definition 1.3]. See discussion at the end of Section 5.
@@ 2 Preliminaries
$\kappa\in(0,8),\qquad a=\frac{2}{\kappa},\qquad h=\frac{6-\kappa}{2\kappa},\qquad c=\frac{(6-\kappa)(3\kappa-8)}{2\kappa}.$ (2.1)
$\mathbb{H}$-hulls and half-plane capacity
@@ 2.1 Pure partition functions
  • The functions satisfy the following power-law bound: there exist constants $C\in(0,\infty)$ and $p\in(0,\infty)$ such that for all $(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$,
    $|\mathcal{Z}_{\alpha}(x_{1},\ldots,x_{2N})|\leq C\prod_{1\leq j<k\leq 2N}|x_{k}-x_{j}|^{\mu_{kj}(p)},\quad\text{where }\mu_{kj}(p)=\begin{cases}p,&\text{if }|x_{k}-x_{j}|\geq 1,\\ -p,&\text{if }|x_{k}-x_{j}|<1.\end{cases}$ (2.3)
The uniqueness when $\kappa\in(0,8)$ of such collection of functions were proved in [FK15a]. The existence is proved in [FK15b, KP16, PW19, Wu20, FLPW24]. Pure partition functions enjoy a refined power law bound which we summarize in the following two lemmas. They will be used later.
Lemma 2.2.
Fix $\kappa\in(0,6]$, pure partition functions satisfy the following refined power law bound:
@@
Lemma 2.3.
Fix $\kappa\in(4,8)$, pure partition functions satisfy the following refined power law bound: there exists constant $C\in(0,\infty)$ (depending only on $N$ and $\kappa$) such that for all $(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$,
$0<\mathcal{Z}_{\alpha}(x_{1},\ldots,x_{2N})\leq C\prod_{1\leq j<k\leq 2N}|x_{k}-x_{j}|^{\mu_{kj}},\quad\text{where }\mu_{kj}=\begin{cases}10/\kappa,&\text{if }|x_{k}-x_{j}|\geq 1,\\ -6/\kappa,&\text{if }|x_{k}-x_{j}|<1.\end{cases}$ (2.5)
The bound (2.4) is only proved for $\kappa\in(0,6]$ (see [Wu20, Theorem 1.7]) and we will prove the bound (2.5) for $\kappa\in(4,8)$ in Section 2.2. Note that the bound (2.5) is stronger than (2.3) but is weaker than (2.4). Nevertheless, it suffices for us to complete the proof of main conclusions.
2.2 Proof of Lemma 2.3
The proof of Lemma 2.3 relies on recent development [FPW24, FLPW24] on the relation between Coulomb gas integrals and pure partition functions, both of which are solutions to BPZ equations (2.2). We first introduce some necessary notations. Fix $\kappa\in(4,8)$. For $N\geq 1$, we define
$\mathcal{G}^{(N)}(x_{1},\ldots,x_{2N}):=\left(\frac{2|\cos(4\pi/\kappa)|\Gamma(2-8/\kappa)}{\Gamma(1-4/\kappa)^{2}}\right)^{N}\int_{x_{1}}^{x_{2}}\cdots\int_{x_{2N-1}}^{x_{2N}}f(\boldsymbol{x};u_{1},\ldots,u_{N})\mathrm{d}u_{1}\cdots\mathrm{d}u_{N},$ (2.6)
where $\mathrm{d}u_{q}$ is integrated from $x_{2q-1}$ to $x_{2q}$ for $1\leq q\leq N$, and the integrand is
$f(\boldsymbol{x};u_{1},\ldots,u_{N}):=\prod_{1\leq j<k\leq 2N}(x_{k}-x_{j})^{2/\kappa}\prod_{1\leq r<s\leq N}|u_{s}-u_{r}|^{8/\kappa}\prod_{1\leq j\leq 2N,1\leq q\leq N}|u_{q}-x_{j}|^{-4/\kappa}.$ (2.7)
The function $\mathcal{G}^{(N)}$ is the Coulomb gas integral function with link pattern
¯=$\boldsymbol{\underline{\cap\cap}}=\{\{1,2\},\ldots,\{2N-1,2N\}\}\in\mathrm{LP}_{N},$
see for instance [FLPW24, Definition 1.4]. We have the following two lemmas for $\mathcal{G}^{(N)}$.
Lemma 2.4.
Fix $\kappa\in(4,8)$. For all $(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$,
$\mathcal{G}^{(N)}(x_{1},\ldots,x_{2N})=\sum_{\alpha\in\mathrm{LP}_{N}}\mathcal{M}_{\nu}(\alpha,\boldsymbol{\underline{\cap\cap}})\mathcal{Z}_{\alpha}(x_{1},\ldots,x_{2N}),$ (2.8)
where $\nu=2|\cos(4\pi/\kappa)|$ and $\mathcal{M}_{\nu}(\alpha,\boldsymbol{\underline{\cap\cap}})=\nu^{\ell(\alpha,\boldsymbol{\underline{\cap\cap}})}$ and $\ell(\alpha,\boldsymbol{\underline{\cap\cap}})$ denotes the number of loops in the meander formed from the two link patterns $\alpha$ and $\boldsymbol{\underline{\cap\cap}}$.
Proof.
See [FPW24, Proposition 1.10] and [FLPW24, Proposition 1.11]. ∎
Lemma 2.5.
Fix $\kappa\in(4,8)$. There exists constant $C\in(0,\infty)$ (depending only on $N$ and $\kappa$) such that for all $(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$,
$0<\mathcal{G}^{(N)}(x_{1},\ldots,x_{2N})\leq C\prod_{1\leq j<k\leq 2N}|x_{k}-x_{j}|^{\mu_{kj}},\quad\text{where }\mu_{kj}=\begin{cases}10/\kappa,&\text{if }|x_{k}-x_{j}|\geq 1,\\ -6/\kappa,&\text{if }|x_{k}-x_{j}|<1.\end{cases}$ (2.9)
Proof.
The second term in the integrand (2.7) can be bounded from above: for $1\leq r<s\leq N$,
$|u_{s}-u_{r}|\leq|x_{2s}-x_{2r-1}|.$
The third term in the integrand (2.7) can be bounded from above: for $1\leq q\leq N$,
$\prod_{1\leq j\leq 2N}|u_{q}-x_{j}|^{-4/\kappa}\leq\prod_{1\leq j<2q-1}|x_{2q-1}-x_{j}|^{-4/\kappa}\times|u_{q}-x_{2q-1}|^{-4/\kappa}|u_{q}-x_{2q}|^{-4/\kappa}\times\prod_{2q<j\leq 2N}|x_{2q}-x_{j}|^{-4/\kappa}.$
Note that, using change of variables $v=\frac{u_{q}-x_{2q-1}}{x_{2q}-x_{2q-1}}$,
$\int_{x_{2q-1}}^{x_{2q}}|u_{q}-x_{2q-1}|^{-4/\kappa}|u_{q}-x_{2q}|^{-4/\kappa}\mathrm{d}u_{q}=$ $(x_{2q}-x_{2q-1})^{1-8/\kappa}\int_{0}^{1}v^{-4/\kappa}(1-v)^{-4/\kappa}\mathrm{d}v$
$=$ $(x_{2q}-x_{2q-1})^{1-8/\kappa}\frac{\Gamma(1-4/\kappa)^{2}}{\Gamma(2-8/\kappa)}.$
Plugging these three observations into (2.6), there exists constant $C\in(0,\infty)$ (depending only on $N$ and $\kappa$),
$\mathcal{G}^{(N)}(x_{1},\ldots,x_{2N})\leq$ $C\prod_{1\leq j<k\leq 2N}(x_{k}-x_{j})^{2/\kappa}\times\prod_{1\leq r<s\leq N}|x_{2s}-x_{2r-1}|^{8/\kappa}$ (2.10)
×$\times\prod_{1\leq q\leq N}\prod_{1\leq j<2q-1}|x_{2q-1}-x_{j}|^{-4/\kappa}\times|x_{2q}-x_{2q-1}|^{1-8/\kappa}\times\prod_{2q<j\leq 2N}|x_{2q}-x_{j}|^{-4/\kappa}.$
In the right-hand side of (2.10),
  • the accumulated exponent for the term $|x_{2q}-x_{2q-1}|$ with $1\leq q\leq N$ is given by
    $2/\kappa+1-8/\kappa=1-6/\kappa;$
  • the accumulated exponent for the term $|x_{2q}-x_{2r-1}|$ with $1\leq r<q\leq N$ is given by
    $2/\kappa+8/\kappa=10/\kappa;$
  • the accumulated exponent for the term $|x_{2q+1}-x_{2q}|$ with $11\leq q\leq N-$ 1 is given by
    $2/\kappa-4/\kappa=-2/\kappa;$
  • the accumulated exponent for the term $|x_{2j+1}-x_{2q}|$ with $11\leq q<j\leq N-$ 1 is given by
    $2/\kappa-4/\kappa-4/\kappa=-6/\kappa;$
  • the accumulated exponent for the term $|x_{2r}-x_{2q}|$ with $1\leq q<r\leq N$ is given by
    $2/\kappa-4/\kappa=-2/\kappa;$
  • the accumulated exponent for the term $|x_{2r-1}-x_{2q-1}|$ with $1\leq q<r\leq N$ is given by
    $2/\kappa-4/\kappa=-2/\kappa.$
Combining these observations, we obtain (2.9) as desired. ∎
Proof of Lemma 2.3.
From (2.8), for any $\alpha\in\mathrm{LP}_{N}$, we have $\mathcal{G}^{(N)}\geq\mathcal{M}_{\nu}(\alpha,\boldsymbol{\underline{\cap\cap}})\mathcal{Z}_{\alpha}$. Combining with (2.9), we obtain (2.5) as desired. ∎

### v3

@@ 1 Introduction
When $N=1$, we denote by $X(\Omega;x_{1},x_{2})$ the set of continuous simple unparameterized curves in $\Omega$ connecting $x_{1}$ and $x_{2}$ such that they only touch the boundary $\partial\Omega$ in $\{x_{1},x_{2}\}$. Chordal SLE in $(\Omega;x_{1},x_{2})$ is a continuous random curve in $\Omega$ connecting $x_{1}$ and $x_{2}$. Note that $\mathrm{SLE}_{\kappa}$ in $(\Omega;x_{1},x_{2})$ is a random curve in $X(\Omega;x_{1},x_{2})$ when $\kappa\leq 4$. In general, for $N\geq 2$ and $\alpha=\{\{a_{1},b_{1}\},\ldots,\{a_{N},b_{N}\}\}\in\mathrm{LP}_{N}$, we denote by $X_{\alpha}(\Omega;x_{1},\ldots,x_{2N})$ the set of families $(\gamma_{1},\ldots,\gamma_{N})$ of curves such that, for each $s\in\{1,\ldots,N\}$, we have $\gamma_{s}\in X(\Omega;x_{a_{s}},x_{b_{s}})$.
@@
Fix $\kappa\in(0,4]$. For $N\geq 1$ and $\alpha\in\mathrm{LP}_{N}$, there exists a unique probability measure on $(\gamma_{1},\ldots,\gamma_{N})\in X_{\alpha}(\Omega;x_{1},\ldots,x_{2N})$ such that, for each $j\in\{1,\ldots,N\}$, the conditional law of $\gamma_{j}$ given $\{\gamma_{1},\ldots,\gamma_{N}\}\setminus\{\gamma_{j}\}$ is the chordal $\mathrm{SLE}_{\kappa}$ connecting $x_{a_{j}}$ and $x_{b_{j}}$ in the connected component of the domain $\Omega\setminus\bigcup_{k\neq j}\gamma_{k}$ having $x_{a_{j}}$ and $x_{b_{j}}$ on its boundary. For the existence and uniqueness of such measures, see [KL07], [PW19, Theorem 1.3] and [BPW21, Theorem 1.2]. We call this measure chordal $N$-$\mathrm{SLE}_{\kappa}$ in polygon $(\Omega;x_{1},\ldots,x_{2N})$ associated to link pattern $\alpha$.
@@
In the companion paper [FWY24], we derived the connection between chordal $N$-$\mathrm{SLE}_{\kappa}$ in the unit disc and Dyson circular ensemble for $\kappa\in(0,4]$. In this article, we will prove corresponding conclusions for chordal $N$-$\mathrm{SLE}_{\kappa}$ and Dyson Brownian motion. The main conclusions are summarized as follows. We will parameterize multiple chordal SLEs by a “common time” and derive the transition density of the driving function in Theorem 1.1: it can be given by the transition density of Dyson Brownian motion and Green’s function (Definition 1.2). Using such connection and the asymptotic of the transition density of Dyson Brownian motion, we derive the asymptotic of the probability of a rare event and show that, conditional on this rare event, the driving function of multiple chordal SLEs converges to Dyson Brownian motion (Proposition 1.3). In particular, we take $\kappa=4$ as an example and connect our result to Gaussian free field level lines (Proposition 4.1). In this case, the driving function has an explicit SDE (4.2), Green’s function has an explicit formula (4.4), and the Dyson Brownian motion is of parameter $\beta=2$. Finally, we give one application of our main result: we derive the asymptotic of the probability for chordal $N$-SLE to hit a small neighborhood of a boundary point (Theorem 1.4).
@@
Let us elaborate on the tools we need to prove our main results. To prove the first main conclusion Theorem 1.1, we first relate global multiple SLEs to multiple SLE partition functions “pure partition functions” (see Section 2.1); second, we use these partition functions to construct $2N$-time local martingales (see Section 3.2); and finally, we take the common time in these local martingales and derive the desired transition density. To prove the second main conclusion Proposition 1.3 and Theorem 1.4, we need properties of pure partition functions and results in complex analysis.
@@ 1.1 Multiple SLEs and Dyson Brownian motion
Fix $\kappa\in(0,4]$ and $\alpha\in\mathrm{LP}_{N}$.
@@
Fix $\kappa\in(0,4]$ and $\alpha\in\mathrm{LP}_{N}$. For $(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$, we define Green’s function in the boundary for chordal $N$-$\mathrm{SLE}_{\kappa}$ as
@@ 1.2 Estimates on multiple SLEs
Fix $\kappa\in(0,4]$, $\alpha\in\mathrm{LP}_{N}$ and $\boldsymbol{x}\in\mathfrak{X}_{2N}$. We have
@@
Fix $\kappa\in(0,4]$, $\boldsymbol{x}=(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$, $\alpha\in\mathrm{LP}_{N}$ and suppose $(\gamma_{1},\ldots,\gamma_{N})\sim\mathbb{P}_{\alpha}^{(\boldsymbol{x})}$. Then we have [^2] For two functions $F_{1}(\boldsymbol{x};R)$ and $F_{2}(\boldsymbol{x};R)$, we write $F_{1}(\boldsymbol{x};R)\asymp F_{2}(\boldsymbol{x};R)$ as $R/|\boldsymbol{x}|\to\infty$ if there exist constants $C\in(1,\infty)$ and $R_{0}$ such that $C^{-1}\leq\frac{F_{1}(\boldsymbol{x};R)}{F_{2}(\boldsymbol{x};R)}\leq C$ for all $R\geq R_{0}|\boldsymbol{x}|$.
@@
The rest of this article is organized as follows. We give preliminaries on multiple SLEs and Dyson Brownian motion in Section 2. We prove Theorem 1.1 in Section 3. The main ingredients for the proof are $2N$-time local martingales constructed in Section 3.2. Using Theorem 1.1 and the asymptotic of the transition density of Dyson Brownian motion, we complete the proof of Proposition 1.3 in Section 3.3. We prove Proposition 4.1 in Section 4. Finally, we prove Proposition 1.4 in Section 5 using Proposition 1.3. The conclusion in Proposition 1.4 is proved for $\kappa\in(0,8),N=1$ in [Law15] and is proved for $\kappa\in(0,8),N=2$ in [Zha21]. We will prove (1.8) in Section 5 for all $\kappa\in(0,4],N\geq 3,\alpha\in\mathrm{LP}_{N}$.
@@ 2 Preliminaries
$\kappa\in(0,4],\qquad a=\frac{2}{\kappa},\qquad h=\frac{6-\kappa}{2\kappa},\qquad c=\frac{(6-\kappa)(3\kappa-8)}{2\kappa}.$ (2.1)
$\mathbb{H}$-hulls and half-plane capacity.
@@ 2.1 Pure partition functions
  • The functions satisfy the following power-law bound:
@@
The uniqueness when $\kappa\in(0,8)$ of such collection of functions were proved in [FK15a]. The existence is proved in [FK15b, KP16, PW19, Wu20, FLPW24].